OpenStudy (anonymous):

Kim works on Social Studies homework for 2 2/5 h. Then she works on Math homework for 1 1/4 h. How many hours total does Kim spend doing homework?

OpenStudy (anonymous):



OpenStudy (solomonzelman):

I will give you an example of a similar problem, and you tell me what to do in your case.

OpenStudy (solomonzelman):

`Question`: Bob and Sam are gathering apples together, to make a jam for a New Year celebration. Bob gathered 46 apples, and Sam gathered 51 apples. How many apples do they have for the New Year celebration altogether? `Solution:` Sam has 46 apples, and Bob has 51 apples, so together they will have ( `46+51` ) apples. `Answer`:They have 97 apples altogether.

OpenStudy (solomonzelman):

Can you tell me what to do in your (similar to my example's) situation?

OpenStudy (anonymous):

can you help me with my problem i dont understand

OpenStudy (solomonzelman):

Kim spent \(\large\color{slate}{ 2\frac{\LARGE 2 }{\LARGE 5} }\) hours on Social Studies, and \(\large\color{slate}{ 1\frac{\LARGE 1 }{\LARGE 4} }\) hours on Math. So how many hours would he spend on `Social Studies` and `Math` together ?

OpenStudy (solomonzelman):

still lost?

OpenStudy (anonymous):

3 3/9

OpenStudy (solomonzelman):

it seems as though you have the answer options that you haven't posted and trying to eliminate the answer until I say "yes".

OpenStudy (solomonzelman):

oh I see what you did

OpenStudy (solomonzelman):

you are correct that you have to add the hours, but you added incorrectly.

OpenStudy (solomonzelman):

\(\large\color{slate}{ 2\frac{\LARGE 2 }{\LARGE 5} +1\frac{\LARGE 1 }{\LARGE 4} }\) this si what you need to calculate.

OpenStudy (solomonzelman):

I will simplify it a bit for you: \(\large\color{slate}{ 2\frac{\LARGE 2 }{\LARGE 5} +1\frac{\LARGE 1 }{\LARGE 4} }\) \(\large\color{slate}{ 2+\frac{\LARGE 2 }{\LARGE 5} +1+\frac{\LARGE 1 }{\LARGE 4} }\) \(\large\color{slate}{ 2+1+\frac{\LARGE 2 }{\LARGE 5} +\frac{\LARGE 1 }{\LARGE 4} }\) \(\large\color{slate}{ 3+\frac{\LARGE 2 }{\LARGE 5} +\frac{\LARGE 1 }{\LARGE 4} }\)

OpenStudy (solomonzelman):

then you need to find the common denominator, not to just add tops and bottoms.

OpenStudy (anonymous):

\[\frac{ a }{ b }+\frac{c} { d }=\frac{ ad+bc }{ bd }\]

OpenStudy (anonymous):

how i find commom denominator

OpenStudy (solomonzelman):

yes, use what Fallontine posted.

OpenStudy (party_girl):

Add them both together to find the answer😃

OpenStudy (anonymous):

what i have to add together

OpenStudy (solomonzelman):

the `hours spent on Social studies` and the `hours spent on Math`

OpenStudy (solomonzelman):

\(\large\color{blue}{ 2\frac{\LARGE 2 }{\LARGE 5} +1\frac{\LARGE 1 }{\LARGE 4} }\) like this.

OpenStudy (solomonzelman):

can you first add, \(\large\color{blue}{ \frac{\LARGE 2 }{\LARGE 5} +\frac{\LARGE 1 }{\LARGE 4} }\) ?

OpenStudy (anonymous):

3/9

OpenStudy (solomonzelman):

oh you are not adding the top and bottoms. this is what you need to do: Steps to add \(\large\color{green}{ \frac{\LARGE 2 }{\LARGE 5} +\frac{\LARGE 1 }{\LARGE 4} }\) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ `1)` multiply the top and the bottom of first fraction by one number and multiply the top and bottom of the second fraction by another number So that the fractions are equivalent. ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ `2)` (after multiplying as directed in step 1) ADD the tops of the fractions together, while leaving the bottom the way it is, and DO NOT add the bottoms together. ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ `3)` if there is a number by which you can divide the top and bottom of a new (formed from addition) fraction, then do so.